OpenStudy (anonymous):

What is the value of x and y? give 2( 2x-8y) + 2 ( 3x-4y ) 4x -16y + 6x -8y 10x -24 y 5x -12y x = 12y/5

OpenStudy (anonymous):

let x=1; then y=5x/12 y=5(1)/12 y=5/12 therefore the value of y when x=1 is 5/12 .......

OpenStudy (anonymous):

Thanks!